# Zhang, Liu, and Zhao

### v3

@@ 10 Details of Convergence Analysis in Section 4
    Under Assumption 1 with $\varepsilon>0$, $\rho>0$, and the corresponding constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$,
@@
Proof 10.9
Proof of Proposition 10.8. For the first statement, since the objective $F$ is a piecewise affine function on a polyhedral set $\Theta$, $\Theta$ can be decomposed into finitely many polyhedral sets such that $F$ is affine on each of them, showing that $r_{d}(\cdot)$ takes finitely many values on $\Theta$. Notice that $\theta\in\mathcal{S}^{d}$ if and only if $r_{d}(\theta)=\max_{d\in\mathcal{T}(\theta;\Theta),\left\lVert d\right\rVert\leq 1}\{-F^{\prime}(\theta;d)\}=0$. Hence, there exists $l_{0}>0$ such that $r_{d}(\theta)\geq l_{0}$ for any $\theta\in\Theta\backslash\mathcal{S}^{d}$. By setting $\eta_{d}\coloneqq\frac{2}{l_{0}}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, we deduce $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq 2\max_{\theta\in\Theta}\left\lVert\theta\right\rVert\leq\eta_{d}r_{d}(\theta)$.
For the second statement, the inequality holds trivially if $\theta\in\mathrm{cl}(\mathcal{S}^{d})$. By Assumption 1, for any $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda_{q}$ and $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(\cdot;\theta^{d},\theta)$ is convex and $\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}$ for all $t\in[0,1].$ According to Drusvyatskiy and Lewis (2018, Theorem 3.3), it follows that
@@
Proposition 10.10
Under Assumption 2, for $\varepsilon\geq 0$ and $\rho>0$, it holds that for any $\theta\in\Theta$, there exists $\bar{\theta}\in\Theta$ such that $\left\lVert\bar{\theta}-\theta\right\rVert\leq 2\,r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\bar{\theta})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}(\theta)$.
A similar result to Proposition 10.10 is provided in Liu et al. (2022, Lemma 4) under the assumption that the relevant surrogation function satisfies Condition (P5). With the surrogation function $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime})$ associated with $r^{\varepsilon,\rho}(\theta^{\prime})$ satisfying (P5), we can prove Proposition 10.10 following the lines of Liu et al. (2022, Lemma 4), and thus the details are omitted here.
Proof 10.11
Proof of Proposition 10.6. For the first statement, let $\{\theta_{t}\}_{t\in\mathcal{K}}$ be a convergent subsequence with the limit point $\theta_{\infty}$. Given $\varepsilon^{\prime}\in[0,\varepsilon)$, it follows from the Lipschitz continuity of $f$ that $\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\subseteq\mathcal{I}^{\varepsilon}(\theta_{t})$ for sufficiently large $t\in\mathcal{K}$. With $\widehat{P}^{\rho}_{I}(\theta_{t})\coloneqq\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta_{t},I)+\frac{\rho}{2}\left\lVert\theta-\theta_{t}\right\rVert^{2}\right\}$, by definition of the residual $r^{\varepsilon,\rho}$ there exists $I^{*}_{t}\in\underset{I\in\mathcal{I}^{\varepsilon}(\theta_{t})}{\arg\min}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon,\rho}(\theta_{t})=\,\big{\lVert}\,{\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}}\,\big{\rVert}$. Then for sufficiently large $t\in\mathcal{K}$, we have that
$\widehat{F}^{*}(\theta_{t})$ $\coloneqq\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
$\geq\min_{I\in\mathcal{I}^{\varepsilon}(\theta_{t})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
$=\widehat{F}(\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t});\theta_{t},I^{*}_{t})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}$
$\geq F(\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t}))+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}.$ (22)
Since $\lim_{t(\in\mathcal{K})\rightarrow\infty}r^{{\varepsilon},{\rho}}(\theta_{t})=0$, we deduce $\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})=\theta_{\infty}$. Additionally, for a fixed mapping $I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})$, the surrogation function $\widehat{F}(\cdot;\cdot,I)$ is continuous on $\Theta\times\Theta$ by assumption, and thus by Shapiro et al. (2021, Theorem 7.23), $\widehat{P}^{\rho}_{I}(\cdot)$ is continuous at $\theta_{\infty}$. It follows from (10.11) that
$F(\theta_{\infty})\leq\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{F}^{*}(\theta_{t})=\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{\infty});\theta_{\infty},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\},$
which leads to $F(\theta_{\infty})<\displaystyle{\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\backslash\mathcal{I}^{0}(\theta_{\infty})}}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{I}(\theta_{\infty});\theta_{\infty},I)+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,\widehat{P}^{\rho^{\prime}}_{I}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\}$ for $\rho^{\prime}>\rho$, yielding $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$.
We next prove the second statement. For the special case of $\varphi$ being piecewise affine, Assumption 2 holds naturally. Given $\varepsilon\geq 0$ and $\rho>0$, by Proposition 10.8 there exists $\eta_{d}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{d}r_{d}(\theta)$ for any $\theta\in\Theta$; by Proposition 10.10, for any $\theta\in\Theta$, there exists $\bar{\theta}\in\Theta$ such that $\left\lVert\bar{\theta}-\theta\right\rVert\leq 2r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\bar{\theta})\leq(2\rho+\frac{5\kappa}{2})r^{\varepsilon,\rho}(\theta)$. Setting $\eta_{r}=2+\eta_{d}\left(2\rho+\frac{5\kappa}{2}\right)$, we have that
$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\eta_{d}r_{d}(\bar{\theta})+2r^{\varepsilon,\rho}(\theta)\leq\eta_{r}r^{\varepsilon,\rho}(\theta).$
To prove the desired statement for general B-differentiable $\varphi$ under Assumptions 2 and 1, we first show that given $\varepsilon>\varepsilon^{\prime}\geq 0$ and $\rho^{\prime}>\rho>0$, for any $\lambda^{\prime}_{d}>0$ there exists $\lambda_{d}>0$ such that
@@
By contradiction, suppose that there exists $\bar{\lambda}^{\prime}_{d}>0$ such that for any $\lambda>0$, there exist points $\theta,\theta^{\prime}\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda$, $\left\lVert\theta-\theta^{\prime}\right\rVert\leq 2\lambda$, and $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime})>\bar{\lambda}^{\prime}_{d}$. Then there exist two sequences $\{\theta_{k}\}_{k}$ and $\{\theta^{\prime}_{k}\}_{k}$ in $\Theta$ with $\inf_{k\in\mathbb{N}}r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime}_{k})\geq\bar{\lambda}^{\prime}_{d}$ such that both $r^{\varepsilon,\rho}(\theta_{k})$ and $\left\lVert\theta_{k}-\theta^{\prime}_{k}\right\rVert$ converge to 0. Since $\Theta$ is compact, by passing to a subsequence if necessary, we can assume that both $\{\theta_{k}\}_{k}$ and $\{\theta^{\prime}_{k}\}_{k}$ converge to $\theta_{\infty}\in\Theta$. By the first statement of the proposition, we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$. Let $I_{k}\in{\arg\min}_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta^{\prime}_{k})}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{I}(\theta^{\prime}_{k});\theta^{\prime}_{k},I)+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{I}(\theta^{\prime}_{k})-\theta^{\prime}_{k}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime}_{k})=\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{I_{k}}(\theta^{\prime}_{k})-\theta^{\prime}_{k}}\,\big{\rVert}>\bar{\lambda}^{\prime}_{d}$. Since the total number of index mappings is finite, by passing to a subsequence if necessary, we can assume that $I_{k}\equiv\widehat{I}$ for any $k\in\mathbb{N}$. By continuity of $\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\cdot)$ at $\theta_{\infty}$, we derive $\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}\geq\bar{\lambda}^{\prime}_{d}$ and $\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty});\theta_{\infty},\widehat{I})+\frac{\rho^{\prime}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}^{2}\leq F(\theta_{\infty})$, implying $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})>0$, a contradiction.
@@
Given $\varepsilon^{\prime}\in(0,\varepsilon)$ and $\rho^{\prime}\in(\rho,+\infty)$, let $\delta_{q}^{\prime},\lambda_{q}^{\prime}$, and $c_{q}^{\prime}$ be the corresponding constants in Assumption 1. There exists $\lambda_{q}>0$ such that (23) holds for $\lambda_{q}^{\prime}$. For $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{\delta_{q}^{\prime}}{3}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{0}\coloneqq\min\left\{\lambda_{q},\frac{\delta_{q}^{\prime}}{3}\right\}$, by Proposition 10.10 let $\bar{\theta}$ be the point satisfying $\left\lVert\bar{\theta}-{\theta}\right\rVert\leq 2\,r^{\varepsilon,\rho}({\theta})\leq 2\lambda_{q}$ and $r_{d}(\bar{\theta})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}({\theta})$. Then we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta})\leq\lambda_{q}^{\prime}$ from (23) and
@@
$\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})\leq\mathrm{dist}(\theta,\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\frac{\delta_{q}^{\prime}}{3}+2r^{\varepsilon,\rho}(\theta)\leq\delta_{q}^{\prime}.$
@@
It follows from Proposition 10.8 that $\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})\leq\frac{2}{c_{q}^{\prime}}r_{d}(\bar{\theta})$. We further obtain that
@@
$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\frac{2}{c_{q}^{\prime}}r_{d}(\bar{\theta})+2r^{\varepsilon,\rho}(\theta)\leq\left(\frac{4\rho+5\kappa}{c_{q}^{\prime}}+2\right)r^{\varepsilon,\rho}(\theta).$
@@
10.4 Verification of Assumption 1
We verify that Assumption 1 holds when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope by proving a stronger result in Claim 2 with the following lemma.
Lemma 10.12 (Cui et al. (2018), Proposition 2 (ii))
@@ 10.4 Verification of Assumption 1
Recall the notations $\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)\coloneqq\mathop{\arg\min}_{\theta^{\prime}\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta^{\prime};\theta)+\frac{\rho}{2}\left\lVert\theta^{\prime}-\theta\right\rVert^{2}\right\}$ and $\bar{F}(t;\theta_{1},\theta_{2})\coloneqq F(\theta_{1}+t(\theta_{2}-\theta_{1}))$ for $\theta_{1},\theta_{2}\in\Theta$. In the analysis below, let $\partial(S)$ denote the boundary of $S$, $\mathrm{int}(S)$ denote the interior of $S$, and $\mathrm{rint}(S)$ denote the relative interior of $S$.
@@
Let $\Theta\subseteq\mathbb{R}^{q}$ be a polytope and $\varphi(\cdot;\xi)$ be a strongly convex function with modulus $c_{\varphi}$ for any $\xi\in S$. Given $\varepsilon>0$ and $\rho>0$, there exist $\lambda_{q}>0$ and $c_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\coloneqq\underset{\theta^{\prime}\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)}{\max}\,\left\lVert\theta-\theta^{\prime}\right\rVert\leq\lambda_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(\cdot;\theta^{d},\theta)$ is convex on $[0,1]$ and
@@
For $\theta\in\mathrm{cl}(\mathcal{S}^{d})$, the claim holds clearly with $\theta^{d}=\theta$. For $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$, it suffices to consider the case with $n=1$, for which we denote $F(\theta)=\varphi(f(\theta))$, where $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ is piecewise linear and $\varphi$ is strongly convex. The polytope $\Theta$ can be partitioned into a finite number of $q$-polytopes, denoted by $\mathfrak{C}\coloneqq\{\mathcal{C}_{i}\}_{i=1}^{N}$, such that $f$ is linear on each of them and the intersection of any two polytopes in $\mathfrak{C}$ is their common face. Given $\mathcal{C}\in\mathfrak{C}$, $f$ restricted to $\mathcal{C}$ can be represented as $f(\theta)=A\theta$ for some $A\in\mathbb{R}^{d\times q}$. Since the strongly convex function $\varphi$ has a unique minimizer on $\{A\theta\in\mathbb{R}^{d}\mid\theta\in\mathcal{C}\}$, the set of minimizers of $F$ on $\mathcal{C}$ denoted by $\mathcal{S}^{*}_{\mathcal{C}}$ can be represented as $\mathcal{S}^{*}_{\mathcal{C}}=[\{\bar{\theta}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ for any $\bar{\theta}\in\mathcal{S}^{*}_{\mathcal{C}}$, where $\mathcal{N}_{A}$ denotes the null space of $A$.

### v4

@@ 8.4 Nonasymptotic convergence result for general cases
Under Assumptions (B1)-(B6), for problem (10)with $\varepsilon>0$, $\rho>0$, and the corresponding constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$ in Assumption (B6),
@@
Proof 8.15
[…]
@@
Proof 8.16
Proof of Proposition 8.13. We first show that given $\varepsilon>\varepsilon^{\prime}\geq 0$ and $\rho^{\prime}>\rho>0$, for any $\lambda^{\prime}_{d}>0$ there exists $\lambda_{d}>0$ such that
@@
By contradiction, suppose that there exists $\bar{\lambda}^{\prime}_{d}>0$ such that for any $\lambda>0$, there exist points $\theta,\bar{\theta}\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda$, $\left\lVert\theta-\bar{\theta}\right\rVert\leq 2\lambda$, and $r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta})>\bar{\lambda}^{\prime}_{d}$. Then there exist two sequences $\{\theta_{k}\}_{k}$ and $\{\bar{\theta}_{k}\}_{k}$ in $\Theta$ with $\inf_{k\in\mathbb{N}}r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta}_{k})\geq\bar{\lambda}^{\prime}_{d}$ such that both $r^{\varepsilon,\rho}(\theta_{k})$ and $\left\lVert\theta_{k}-\bar{\theta}_{k}\right\rVert$ converge to 0. Since $\Theta$ is compact, by passing to a subsequence if necessary, we can assume that both $\{\theta_{k}\}_{k}$ and $\{\bar{\theta}_{k}\}_{k}$ converge to $\theta_{\infty}\in\Theta$. By the first statement of the proposition, we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$. Let $\mathbf{I}_{k}\in{\arg\min}_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\bar{\theta}_{k})}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\bar{\theta}_{k});\bar{\theta}_{k},\mathbf{I})+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\bar{\theta}_{k})-\bar{\theta}_{k}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta}_{k})=\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\mathbf{I}_{k}}(\bar{\theta}_{k})-\bar{\theta}_{k}}\,\big{\rVert}>\bar{\lambda}^{\prime}_{d}$. Since the total number of index combination is finite, by passing to a subsequence if necessary, we can assume that $\mathbf{I}_{k}\equiv\widehat{\mathbf{I}}$ for any $k\in\mathbb{N}$. By continuity of $\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\cdot)$ at $\theta_{\infty}$, we derive $\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}\geq\bar{\lambda}^{\prime}_{d}$ and $\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\theta_{\infty});\theta_{\infty},\widehat{\mathbf{I}})+\frac{\rho^{\prime}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}^{2}\leq F(\theta_{\infty})$, implying $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})>0$, a contradiction.
@@
Given $\varepsilon^{\prime}\in(0,\varepsilon)$ and $\rho^{\prime}\in(\rho,+\infty)$, let $\delta_{q}^{\prime},\lambda_{q}^{\prime}$, and $c_{q}^{\prime}$ be the corresponding constants in Assumption (B6). There exists $\lambda_{q}>0$ such that (31) holds for $\lambda_{q}^{\prime}$. For $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{\delta_{q}^{\prime}}{3}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{0}\coloneqq\min\left\{\lambda_{q},\frac{\delta_{q}^{\prime}}{3}\right\}$, by Proposition 8.10 let $\theta^{\prime}$ be the point satisfying $\left\lVert\theta^{\prime}-{\theta}\right\rVert\leq 2\,r^{\varepsilon,\rho}({\theta})\leq 2\lambda_{q}$ and $r_{d}(\theta^{\prime})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}({\theta})$. Then we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime})\leq\lambda_{q}^{\prime}$ from (31) and
@@
$\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})\leq\mathrm{dist}(\theta,\mathcal{S}^{d})+\left\lVert\theta-\theta^{\prime}\right\rVert\leq\frac{\delta_{q}^{\prime}}{3}+2r^{\varepsilon,\rho}(\theta)\leq\delta_{q}^{\prime}.$
@@
It follows from Proposition 8.8 that $\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})\leq\frac{2}{c_{q}^{\prime}}r_{d}(\theta^{\prime})$. We further obtain that
@@
$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})+\left\lVert\theta-\theta^{\prime}\right\rVert\leq\frac{2}{c_{q}^{\prime}}r_{d}(\theta^{\prime})+2r^{\varepsilon,\rho}(\theta)\leq\left(\frac{4\rho+5\kappa}{c_{q}^{\prime}}+2\right)r^{\varepsilon,\rho}(\theta).$
@@
8.5 Verification of Assumption (B6)
We verify that Assumption (B6) holds when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope by proving a stronger result in Claim 1 with the following lemma.
[…]
@@ 8.5 Verification of Assumption (B6)
Recall the notations $\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)\coloneqq\mathop{\arg\min}_{\theta^{\prime}\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta^{\prime};\theta)+\frac{\rho}{2}\left\lVert\theta^{\prime}-\theta\right\rVert^{2}\right\}$ and $\bar{F}(t;\theta_{1},\theta_{2})\coloneqq F(\theta_{1}+t(\theta_{2}-\theta_{1}))$ for $\theta_{1},\theta_{2}\in\Theta$. In the analysis below, let $\partial(S)$ denote the boundary of $\Xi$, $\mathrm{int}(S)$ denote the interior of $\Xi$, and $\mathrm{rint}(S)$ denote the relative interior of $\Xi$.
@@
Let $\Theta\subseteq\mathbb{R}^{q}$ be a polytope and $\varphi(\cdot,\xi)$ be a strongly convex function with modulus $c_{\varphi}$ for any $\xi\in\Xi$. Given $\varepsilon>0$ and $\rho>0$, there exist $\lambda_{q}>0$ and $c_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\coloneqq\underset{\theta^{\prime}\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)}{\max}\,\left\lVert\theta-\theta^{\prime}\right\rVert\leq\lambda_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(\cdot;\theta^{d},\theta)$ is convex on $[0,1]$ and
@@
For $\theta\in\mathrm{cl}(\mathcal{S}^{d})$, the claim holds clearly with $\theta^{d}=\theta$. For $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$, it suffices to consider the case with $n=1$, for which we denote $F(\theta)=\varphi(f(\theta))$, where $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ is piecewise linear and $\varphi$ is strongly convex. The polytope $\Theta$ can be partitioned into a finite number of $q$-polytopes, denoted by $\mathfrak{C}\coloneqq\{\mathcal{C}_{i}\}_{i=1}^{N}$, such that $f$ is linear on each of them and the intersection of any two polytopes in $\mathfrak{C}$ is their common face. Given $\mathcal{C}\in\mathfrak{C}$, $f$ restricted to $\mathcal{C}$ can be represented as $f(\theta)=A\theta$ for some $A\in\mathbb{R}^{d\times q}$. Since the strongly convex function $\varphi$ has a unique minimizer on $\{A\theta\in\mathbb{R}^{d}\mid\theta\in\mathcal{C}\}$, the set of minimizers of $F$ on $\mathcal{C}$ denoted by $\mathcal{S}^{*}_{\mathcal{C}}$ can be represented as $\mathcal{S}^{*}_{\mathcal{C}}=[\{\theta^{\prime}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ for any $\theta^{\prime}\in\mathcal{S}^{*}_{\mathcal{C}}$, where $\mathcal{N}_{A}$ denotes the null space of $A$.